SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: March 03, 2008
Shift Taker: David Doll
-----------
General Features:
(theta,phi)

Warm crystals in hitmap	(52,63), (56,63), (51,64)
Warm crystals in 0.01 GeV hitmap (17,43), (30,45), (32,46), (36,64)
Warm crystals in 0.1 GeV hitmap (30,45), (32,46), (36,64)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.1 GeV hitmap (29,84) 
Cool crystal  in 1 GeV hitmap    		

[1] =  TC Damage:OOS/OOS ~10^-3 in 1-3 and 1-6 all fibers.

0080122:[1]
0080129: 2 noisy clusters outside of the digi timing peak window.
                A slight second peak in the no cuts digi multiplicity around 2000.
                TC Damage:OOO and OOS ~10^-3 in 0-4, all fibers.
0080130: Additional warm crystal in hitmap at (36,48).
                TC Damage:OOO and OOS ~10^-3 in 0-4, all fibers.
0080131: Additional warm crystal in hitmap at (36,48).
                Additional warm crystals in E=0.1GeV at (30,45) and (32,46).
0080132: Additional warm crystal in hitmap at (36,48).
                Additional warm crystals in E=0.1GeV at (30,45) and (32,46).
                A slight second peak in the no cuts digi multiplicity around 2500 and 3500
                TC Damage:OOO and OOS ~10^-3 in 0-4, all fibers.
0080133: Additional warm crystal in hitmap at (36,48).
                A slight second peak in the no cuts digi multiplicity around 2000.
                TC Damage:OOO and OOS ~10^-3 in 0-4, all fibers.
0080134: Additional warm crystal in hitmap at (36,48).
                Additional warm crystals in E=0.1GeV at (30,45) and (32,46).
                A slight second peak in the no cuts digi multiplicity around 2000.
                TC Damage:OOO and OOS ~10^-3 in 0-4, all fibers.
0080135: A sharp spike in the D2starplus digi timing plot.
                Additional warm crystal in hitmap at (36,48).
                Additional warm crystals in E=0.1GeV at (30,45) and (32,46).
0080136: A sharp spike in the D2starplus digi timing plot.
                Additional warm crystal in hitmap at (36,48). 
                Additional warm crystals in E=0.1GeV at (30,45) and (32,46). 
                TC Damage:OOO and OOS ~10^-3 in 0-4, all fibers.
0080137: Cosmics, everything's fine.
0080139: Cosmics, everything's fine.
0080140: Cosmics, everything's fine.
0080141: Cosmics, everything's fine.
0080142: Cosmics, everything's fine.
0080143: Cosmics, everything's fine.
0080144: Cosmics, everything's fine.
0080145: Cosmics, everything's fine.
0080146: Cosmics, everything's fine.
0080147: Cosmics, everything's fine.
0080148: Incorrectly stored file, only 4 pages!
0080149: Incorrectly stored file, only 4 pages!
0080150: Cosmics, everything's fine.
0080151: Cosmics, everything's fine.


Crystal Spectra:
Badness 1 theta (37-40, 43).

Lightpulser:
4 channels dead, 1 channels out of time. Last updated on: 03-March 2008 1:10 AM PST


 

  Links

  • Shift Report Index
  • Back to the Shifter instructions